# An algebraic-geometric construction of “lump” solutions of the KP1 equation

### v1

@@ Abstract.
In this note, we show how certain everywhere-regular real rational function solutions of the KP1 equation (“multi-lumps”) can be constructed via the polynomial analogs of theta functions from singular rational curves with cusps. The method we use can be understood as producing a degeneration of the well-understood soliton solutions from nodal singular curves. Hence it can be seen as a variation on the long-wave limit technique of Ablowitz and Satsuma from [1, 12], as developed by Zhang, Yang, Li, Guo, and Stepanyants, [13]. We present an explicit example of a three-lump solution constructed via the polynomial analog of the theta function from a rational curve with two cuspidal singular points, each with semigroup $\langle 2,5\rangle$. (In the theory of curve singularities, these are known as $A_{4}$ double points.) We conjecture that these ideas will generalize to give similar $M$-lump solutions with $M=\frac{N(N+1)}{2}$ for $N>2$ starting from rational curves with two singular points with semigroup $\langle 2,2N+1\rangle$ ($A_{2N}$ double points). Similar solutions have been constructed by other methods previously; our contribution is to show how they arise from the algebraic-geometric setting by considering singular curves with several cusps, as in [2].
@@ 1. Introduction
Various classes of solutions to KP1 and KP2 (and other related soliton equations) using techniques from algebraic geometry have been known since the late 1970’s, based on the so-called Krichever construction, which shows how to produce points of the Sato Grassmannian starting from data from an algebraic curve. As a result, the KP equation has generated a great deal of interest in algebraic geometry. In the most spectacular connection, KP solutions play a key role in the results of Shiota, Mulase, and many others on the Schottky problem of characterizing the Jacobian varieties of smooth curves among all principally polarized abelian varieties. In the algebraic-geometric context, the solutions produced from smooth curves via theta functions on their Jacobians have received the greatest amount of attention. In the PDE context, these are the so-called quasiperiodic solutions.
However, it has long been understood–and the details have become increasingly clear–that other classes of KP solutions have connections with other classes of singular curves in a very parallel manner. For instance, one of the classes of solutions that will be very important for us here are the soliton solutions. These arise in the algebraic-geometric context by considering the limit of the theta function as a family of smooth curves of genus $g$ degenerates to an irreducible rational curve with $g$ ordinary double points (“nodes”). The connections here were glimpsed very early and exploited by Mumford for the construction of KdV and KP solitons in [10]. They have been studied in much more detail recently in the works of many authors. The most relevant for our purposes are the papers by Agostini, Fevola, Mandelshtam, and Sturmfels, [3], and the related work of Fevola and Mandelshtam, [7]. The fact that all real regular KP soliton solutions corresponding to $\tau$-functions from points of the totally nonnegative Grassmannian can be expressed by the theta functions on nodal singular curves has recently been established by Kodama in [8].
The connection between rational function KP solutions, Schur polynomials and polynomial analogs of theta functions from cuspidal singular curves also has a long history. We mention in particular the foundational article of Buchstaber, Leykin, and Enolski, [4]. More recently, this connection been studied in the article [2], which gives more details about the relation between the polynomial analogs of theta functions for cuspidal curves and rational KP solutions. We will make use of several crucial results from that article. We will also follow many of the notational and terminological conventions for singular curves established there, so readers may wish to consult [2] for background material.
@@
In this connection, we begin by mentioning the early work of Ablowitz and Satsuma, [1, 12], which showed how to produce regular rational solutions from solitons by what they called the “long-wave limit” process. In our terms, it can be seen easily that their method amounts to taking the theta function from an irreducible rational nodal curve and determining what happens when the nodes degenerate to ordinary cusps (i.e. singular points analytically isomorphic to the origin on $y^{2}-x^{3}=0$, or so-called $A_{2}$ double points). In our terms, the nodes come by identifying pairs of points $\{b,c\}$ on the normalization. We always assume our curves are irreducible and rational, so this is just the complex projective line $\mathbb{P}^{1}$ (i.e. the Riemann sphere). Then Ablowitz and Satsuma’s construction amounts to taking a limit as the pairs $\{b,c\}$ coalesce to single points. Special choices must then be made to ensure that the limit of the soliton solution is regular, essentially by making the limiting polynomial $\tau$-function expressible as a sum of squares of real polynomials with a positive nonzero constant term. Those conditions have also been realized by using other methods to find the limiting solutions, most notably the Gram matrix techniques used by Chakravarty and Zowada in [5, 6], and the perturbation processes described by Zhang, et al. in [13].
In this note, we will produce an example of a regular three-lump rational solution of KP1 by using the constructions from [2] and the idea of degenerating a nodal curve to a cuspidal curve (hence degenerating a soliton solution to a rational function solution) in a particular well-chosen way. The idea is to start from a certain soliton solution coming from the theta function of an irreducible nodal curve of arithmetic genus $g=4$, obtained by identifying four pairs of points $\{b_{i},c_{i}\}$, $i=1,\cdots,4$, on $\mathbb{P}^{1}$. By letting $\{b_{1},c_{1},b_{2},c_{2}\}$ coalesce in a particular way, we produce a singular point with semigroup $\langle 2,5\rangle$ (an $A_{4}$ double point). Simultaneously, $\{b_{3},c_{3},b_{4},c_{4}\}$ coalesce to a second, distinct but analytically equivalent singular point, also with semigroup $\langle 2,5\rangle$. From [2], we know that the total degree of the polynomial theta function will be $6=3+3$ in this case because the Young diagram for each of the singular points is the triangular diagram corresponding to the partition $3=2+1$. The connection with these particular Young diagrams and the necessity of degenerating to a curve with two cuspidal singular points was suggested by the results of [5, 6].
@@
From the results of [2], we have a precise recipe for producing the point of the Sato Grassmannian corresponding to the cuspidal curve of arithmetic genus $g=4$. We also know how the corresponding $\tau$-function for the KP solution is related to the polynomial analog of the theta function for the cuspidal curve. To be clear, we note that the results of [2] are geared toward producing solutions of the KP2 equation. Hence we must apply the complex rescaling $y\mapsto i\cdot y$ to get a KP1 solution. We then conclude by finding a $\tau$-function, hence a solution $u(x,y,t)$, that is real for all real $(x,y,t)$. We also show that our solution is regular for all real $(x,y,t)$ by expressing the $\tau$-function as a sum of squares of real polynomials with a nonzero constant term. Throughout this work, we will report results whose derivations made rather heavy use of symbolic computation. We used the Maple 2024 computer algebra system, [9], to work with some of the rather complicated formulas involved.

### v2

@@ Abstract.
In this note, we show how certain everywhere-regular real rational function solutions of the KP1 equation (“multi-lumps”) can be constructed via the polynomial analogs of theta-functions from singular rational curves with cusps. The method we use can be understood either directly, or as producing a degeneration of the well-understood soliton solutions from nodal singular curves. From the second of these points of view, it can be seen as a variation on the long-wave limit technique of Ablowitz and Satsuma from [1, 13], as developed by Zhang, Yang, Li, Guo, and Stepanyants, [14]. Our main explicit example is a three-lump solution constructed via the polynomial analog of the theta-function from a rational curve with two cuspidal singular points, each with semigroup $\langle 2,5\rangle$. (In the theory of curve singularities, these are known as $A_{4}$ double points.) We show that there are similar six-lump solutions from a rational curve with two cusps, each with semigroup $\langle 2,7\rangle$ ($A_{6}$ double points). We conjecture that these ideas will generalize to give similar $M$-lump solutions with $M=\frac{N(N+1)}{2}$ for $N>2$ starting from rational curves with two singular points with semigroup $\langle 2,2N+1\rangle$ ($A_{2N}$ double points). Finally we present a five-lump solution constructed from a rational curve with two cusps, each with semigroup $\langle 3,4\rangle$. Similar solutions have been constructed by other methods previously; our contribution is to show how they arise from the algebraic-geometric setting by considering singular curves with several cusps, as in [2].
@@ 1. Introduction
Various classes of solutions to KP1 and KP2 (and other related soliton equations) using techniques from algebraic geometry have been known since the late 1970’s, based on the so-called Krichever construction, which shows how to produce points of the Sato Grassmannian starting from data from an algebraic curve. As a result, the KP equation has generated a great deal of interest in algebraic geometry. In the most spectacular connection, KP solutions play a key role in the results of Shiota, Mulase, and many others on the Schottky problem of characterizing the Jacobian varieties of smooth curves among all principally polarized abelian varieties. In the algebraic-geometric context, the solutions produced from smooth curves via theta-functions on their Jacobians have received the greatest amount of attention. In the PDE context, these are the so-called quasiperiodic solutions.
However, it has long been understood–and the details have become increasingly clear–that other classes of KP solutions have connections with other classes of singular curves in a very parallel manner. For instance, one of the classes of solutions that will enter for us at one point are the soliton solutions. These arise in the algebraic-geometric context by considering the limit of the theta-function as a family of smooth curves of genus $g$ degenerates to an irreducible rational curve with $g$ ordinary double points (“nodes”). The connections here were glimpsed very early and exploited by Mumford for the construction of KdV and KP solitons in [10]. They have been studied in much more detail recently in the works of many authors. The most relevant for our purposes are the papers by Agostini, Fevola, Mandelshtam, and Sturmfels, [3], and the related work of Fevola and Mandelshtam, [7]. The fact that all real regular KP soliton solutions corresponding to $\tau$-functions from points of the totally nonnegative Grassmannian can be expressed by the theta-functions on nodal singular curves has recently been established by Kodama in [8].
The connection between rational function KP solutions, Schur polynomials and polynomial analogs of theta-functions from cuspidal singular curves also has a long history. We mention in particular the examples from Pelinovskii and Stepanyants, [11], and the foundational article of Buchstaber, Leykin, and Enolski, [4]. More recently, this connection been studied in the article [2], which gives more details about the relation between the polynomial analogs of theta-functions for cuspidal curves and rational KP solutions. We will make use of several crucial results from that article. We will also follow many of the notational and terminological conventions for singular curves established there, so readers may wish to consult [2] for background material.
@@
In this note, we will produce several examples of regular multi-lump rational solutions of KP1 by using the constructions from [2]. Our first example will be a three-lump solution. A similar computation yields a six-lump solution. In both of these, outside a central range of $t$-values, the lumps appear in triangular configurations and undergo an interesting interaction in that central $t$-range. We conjecture that the constructions used for these will generalize to give triangular $N(N+1)/2$-lump solutions for all $N\geq 1$. Finally, we will present a somewhat similar five-lump solution obtained by the same methods.
@@
From the results of [2], we have a precise recipe for producing a point of the Sato Grassmannian corresponding to a cuspidal curve. We also know how the corresponding $\tau$-function for the KP solution is related to the polynomial analog of the theta-function for the cuspidal curve. To be clear, we note that the results of [2] are geared toward producing solutions of the KP2 equation. Hence we must apply the complex rescaling $y\mapsto i\cdot y$ to get a KP1 solution. We then conclude by finding a $\tau$-function, hence a solution $u(x,y,t)$, that is real for all real $(x,y,t)$. We also show that our solution is regular for all real $(x,y,t)$ by expressing the $\tau$-function as a sum of squares of real polynomials with a nonzero constant term.
We will see that the $\tau$-function be computed directly, or by using the idea of degenerating a nodal curve to a cuspidal curve (hence degenerating a soliton solution to a rational function solution) in a particular well-chosen way. In this connection, we begin by mentioning the early work of Ablowitz and Satsuma, [1, 13], which showed how to produce regular rational solutions from solitons by what they called the “long-wave limit” process. In our terms, it can be seen easily that their method amounts to taking the theta-function from an irreducible rational nodal curve and determining what happens when the nodes degenerate to ordinary cusps (i.e. singular points analytically isomorphic to the origin on $y^{2}-x^{3}=0$, or so-called $A_{2}$ double points). In our terms, the nodes come by identifying pairs of points $\{b,c\}$ on the normalization. We always assume our curves are irreducible and rational, so this is just the complex projective line $\mathbb{P}^{1}$ (i.e. the Riemann sphere). Then Ablowitz and Satsuma’s construction amounts to taking a limit as the pairs $\{b,c\}$ coalesce to single points. Special choices must then be made to ensure that the limit of the soliton solution is regular, essentially by making the limiting polynomial $\tau$-function expressible as a sum of squares of real polynomials with a positive nonzero constant term. Those conditions have also been realized by using other methods to find the limiting solutions, most notably the Gram matrix techniques used by Chakravarty and Zowada in [5, 6], and the perturbation processes described by Zhang, et al. in [14].
In our three-lump example, the soliton solution comes from the theta-function of an irreducible nodal curve of arithmetic genus $g=4$, obtained by identifying four pairs of points $\{b_{i},c_{i}\}$, $i=1,\cdots,4$, on $\mathbb{P}^{1}$. By letting $\{b_{1},c_{1},b_{2},c_{2}\}$ coalesce in a particular way, we produce a singular point with semigroup $\langle 2,5\rangle$ (an $A_{4}$ double point). Simultaneously, $\{b_{3},c_{3},b_{4},c_{4}\}$ coalesce to a second, distinct but analytically equivalent singular point, also with semigroup $\langle 2,5\rangle$. From [2], we know that the total degree of the polynomial theta-function will be $6=3+3$ in this case because the Young diagram for each of the singular points is the triangular diagram corresponding to the partition $(2,1)$. The connection with these particular Young diagrams and the necessity of degenerating to a curve with two cuspidal singular points was suggested by the results of [5, 6].
Throughout this work, we will report results whose derivations made rather heavy use of symbolic computation. We used the Maple 2024 computer algebra system, [9], to work with some of the rather complicated formulas involved.
@@
2. Background on cuspidal singular curves and the polynomial analogs of theta-functions
We begin by recalling several results from [2] that will be the basis for a direct method for constructing lump solutions. See the references there for general background on singular curves.
Let $C$ be a reduced, irreducible rational curve of arithmetic genus $n$, Let $\omega_{C}$ be the sheaf of dualizing differentials (an invertible sheaf if $C$ is Gorenstein). The global sections of $\omega_{C}$ can be described as follows. Consider the normalization $\nu:\mathbb{P}^{1}\to C$. The elements of $H^{0}(C,\omega_{C})$ are meromorphic differentials $\omega=g(v)\ dv$ on $\mathbb{P}^{1}$ with
(3) $\sum_{q\in\nu^{-1}(p)}{\rm Res}_{q}(\nu^{*}(f)\cdot\omega)=0$
for all $p\in C$ and all $f\in{\mathcal{O}}_{p}$. If $p$ is unibranch, and $q=\nu^{-1}(p)$ then we must have ${\rm Res}_{q}(\omega)=0$.
The generalized Jacobian of $C$ is defined as the quotient $J(C)=H^{0}(C,\omega_{C})^{*}/H_{1}(C,\mathbb{Z})$. Let $C_{0}\subset C$ be the smooth locus. Fix $P_{0}\in C_{0}$. The Abel mapping with basepoint $P_{0}$ is $\alpha:C_{0}\to J(C)$ defined by $P\mapsto(\omega\mapsto\int_{P_{0}}^{P}\omega)$. This extends by linearity to a mapping on effective divisors of all degrees $d$: $\alpha^{(d)}:C_{0}^{(d)}\to J(C)$. If (arithmetic) genus $g(C)=n$, the subvariety $W_{n-1}$, that is, the closure of the image of $\alpha^{(n-1)}$, will be called the theta-divisor of $C$ and we will consider an implicit equation for the theta-divisor as an analog of a theta-function for $C$. The generalized Jacobian $J(C)\cong\mathbb{C}^{n}$ and the theta-divisor (considered as a hypersurface in $\mathbb{C}^{n}$) is defined by a polynomial equation if and only if $C$ is rational and all singular points of $C$ are “cuspidal” (that is, analytically irreducible, or unibranch).
In [2], a number of general facts concerning the form of the polynomial analogs of theta-functions were derived for cuspidal curves. These depend on two different invariants of the cuspidal singularities. The first of these is the semigroup of the singular point $P$, the set of non-negative integers that are vanishing orders of elements of the local ring ${\mathcal{O}}_{P}$ at $P$. This a subsemigroup of $\mathbb{N}_{\geq 0}$ under addition, and its complement in $\mathbb{N}_{\geq 0}$ is finite.
Theorem 1 ([2], Theorem 1.1).
If $C$ is Gorenstein, rational, of genus $n$ with only cuspidal singularities, the degree of the polynomial analog of the theta-function giving an implicit equation of $W_{n-1}$ is $\leq n(n+1)/2$. The bound is attained for curves with a single singular point with semigroup $\langle 2,2n+1\rangle$.
The complement of the semigroup is known as the set of gaps and the number of gaps is denoted by $\delta_{P}$. When a curve has several singular points, its arithmetic genus is the sum $\sum_{P\in C\setminus C_{0}}\delta_{P}$.
Since the set of gaps is finite, there will be a minimal integer $d_{P}$ such that all $n\geq d_{P}$ are contained in the semigroup. The Gorenstein condition is equivalent to the equation $d_{P}=2\delta_{P}$ at all singular points. We will only consider singularities satisfying this condition in this note.
If the gaps at $P$ are $w_{1}<w_{2}<\cdots<w_{n}$, then the second invariant of the singularity that we will use is the Weierstrass partition $\lambda_{P}=(\lambda_{1},\ldots,\lambda_{n})$ associated to $P$. The parts of $\lambda_{P}$ are defined by
(4) $\lambda_{i}=w_{n+1-i}-(n-i),\text{ for }1\leq i\leq n.$
Example. Consider the rational monomial curve $C$ parametrized by
$\nu:\mathbb{P}^{1}$ $\longrightarrow\mathbb{P}^{3}$
$(t:s)$ $\longmapsto(x_{0}:x_{1}:x_{2}:x_{3})=(t^{6}:t^{2}s^{4}:ts^{5}:s^{6}).$
$C$ is a curve of degree $6$ and arithmetic genus $4$ with implicit equations
$x_{1}x_{3}-x_{2}^{2}=x_{1}^{3}-x_{0}x_{3}^{2}=0.$
It has exactly one singular point $P=\nu(1:0)=(1:0:0:0)$. Using the affine coordinate $v=t/s$ on $\mathbb{P}^{1}$, the singularity is at $v=\infty$ and $P_{0}$, defined by $v=0$, is a smooth point. The local ring ${\mathcal{O}}_{P}$ generated by $(s/r)^{4},(s/t)^{5},(s/t)^{6}$, or $v^{-4},v^{-5},v^{-6}$. That singular point is clearly unibranch with semigroup $=\langle 4,5,6\rangle$. The gaps are $\{1,2,3,7\}$. Note that $\delta=4,d=8$, so this is Gorenstein. By (4), the partition corresponding to the singularity at $P$ is $\lambda=(4,1,1,1)$. By (3), a basis for $H^{0}(C,\omega_{C})$: is $\{dv,\ v\ dv,\ v^{2}\ dv,\ v^{6}\ dv\}$. The theta-divisor is parametrized by
[…]
These abelian integrals are also polynomials, so the 3-fold parametrized above has a polynomial equation. We can find it, for instance, using Gröbner bases. Carrying out the integration, then eliminating $t_{1},t_{2},t_{3}$, we get the implicit equation
(5) $Z_{1}^{7}+21Z_{1}^{4}Z_{3}-84Z_{1}^{3}Z_{2}^{2}+252Z_{1}Z_{3}^{2}+252Z_{2}^{2}Z_{3}-252Z_{4}=0.$
This is the polynomial analog of the theta-function in this case and the pattern described in Theorem 1 above is clear since $7<(4\cdot 5)/2=10$. $\triangle$
In general, the total degree of the polynomial analog of the theta-function for a Gorenstein cuspidal curve and the form of its highest degree term are described as follows.
Theorem 2 ([2], Theorem 5.3).
Let $C$ be a curve with several Gorenstein cuspidal singularities $P_{1},\ldots,P_{h}$ with corresponding partitions $\lambda(P_{1}),\ldots,\lambda(P_{h})$. Take a basis of $H^{0}(C,\omega_{C})$ that is a union of the local bases at the $P_{i}$. For each $i$, let $Z_{1}^{(i)}$ correspond to the basis differential with minimal pole order at $P_{i}$. Then the highest-degree monomial in the polynomial analog of the theta-function is
$(Z_{1}^{(1)})^{|\lambda(P_{1})|}\cdots(Z_{1}^{(h)})^{|\lambda(P_{h})|},$
total degree is $\deg(\theta)=\sum_{i=1}^{h}|\lambda(P_{i})|$.
In the foregoing example, note $\lambda=(4,1,1,1)$. Hence $|\lambda|=7$ and the highest degree term in (5) is $Z_{1}^{7}$. The polynomial there is also homogeneous for the weighting
${\rm wt}(Z_{1})=1,{\rm wt}(Z_{2})=2,{\rm wt}(Z_{3})=3,{\rm wt}(Z_{4})=7.$
These weights are the gaps of the semigroup $\langle 4,5,6\rangle$. This homogeneity is a special feature coming from the $\mathbb{C}^{\ast}$ torus action on the monomial curve.
3. From the theta-function to a KP tau-function
There is a close connection between the polynomial analogs of theta-functions for cuspidal curves as developed in the previous section and tau-functions for the KP hierarchy. The underlying reason for this is that the Krichever construction producing a point of the Sato Grassmannian from data connected to an algebraic curve also applies in this situation of cuspidal singular curves. See [2], Proposition 3.1 for the exact statement. In particular, there is a frame for the subspace corresponding to the Grassmannian point where the columns at the right of the frame matrix come from the series expansions of the differentials in a basis of $H^{0}(C,\omega_{C})$. This yields the following statement.
Theorem 3 ((ACL), Theorem 3.5, Theorem 4.11).
Let $C$ be a Gorenstein cuspidal curve of arithmetic genus $g$. Let $\omega_{j}=\sum_{i=0}^{\infty}a_{ij}u^{i}$ be the local expansions of the differentials in a basis of $H^{0}(C,\omega_{C})$ at the smooth point $P_{0}$, and let $A$ be the $g\times\infty$ transposed matrix $A=(a_{ji})$. Let ${\bf t}=(t_{1},t_{2},\ldots\ )$ be an infinite collection of indeterminates. Then a KP tau-function corresponding to the point of the Sato Grassmannian obtained from $C$ by the Krichever construction is given by
$\tau({\bf t})=e^{\ell({\bf t})}\theta(A{\bf t}+{\bf b}),$
where $\ell$ is a linear function of ${\bf t}$ and ${\bf b}$ is a constant vector.
4. The three-lump example
Consider a cuspidal singular curve with two cusps with semigroup $\langle 2,5\rangle$ at the points $z=\pm 1$ in one affine coordinate on $\mathbb{P}^{1}$. By the constructions from [2], the $\langle 2,5\rangle$-cusps correspond to the triangular Young diagrams from the partition $\lambda=(2,1)$. Moreover, by Theorem 2 the total degree of the polynomial analog of the theta-function will be $3+3=6$ in this case. We take a basis of $H^{0}(C,\omega_{C})$ as follows
$\omega_{1}$ $=\frac{4}{(u-1)^{2}}\ du$
(6) $\omega_{2}$ $=\frac{-12u^{2}}{(u-1)^{4}}\ du$
$\omega_{3}$ $=\frac{4}{(u+1)^{2}}\ du$
$\omega_{4}$ $=\frac{-12u^{2}}{(u+1)^{4}}\ du.$
(The constant factors $4,-12$ are not important here; they resulted from a different calculation to be described later.)
We can find an implicit equation of the $W_{3}$ subvariety of the generalized Jacobian of the cuspidal curve starting from the usual abelian integral parametrization. We take the base point of the abelian integrals as the point $z=0$ or $u=\infty$ on $\mathbb{P}^{1}$, the normalization of the cuspidal curve:
(7) $Z_{j}=\int_{\infty}^{t_{1}}\omega_{j}+\int_{\infty}^{t_{2}}\omega_{j}+\int_{\infty}^{t_{3}}\omega_{j},\quad j=1,\ldots,4.$
Eliminating the $t_{i}$ via a Gröbner basis calculation yields an implicit equation involving a polynomial analog of the theta-function:
$-32256Z_{3}-5376Z_{4}=0.$
As expected by Theorem 2, the highest-degree term is the $Z_{1}^{3}Z_{3}^{3}$ of total degree $6$.
From Theorem 3 (Theorem 4.11 of [2]), to produce a KP tau-function from this theta-function (up to an exponential factor that does not contribute anything when we apply (2) to produce the actual KP1 solution $u(x,y,t)$), we need to compute a frame for the point of the Sato Grassmannian corresponding to the cuspidal curve. In fact, we have already done the relevant computations needed here, since the crucial part of this comes from the series expansions of the dualizing differentials from (4) above. We have
$\omega_{1}$ $=(4+8u+12u^{2}+16u^{3}+\cdots\ )\ du$
(9) $\omega_{2}$ $=(-12u^{2}-48u^{3}+\cdots\ )\ du$
$\omega_{3}$ $=(4-8u+12u^{2}-16u^{3}+\cdots\ )\ du$
$\omega_{4}$ $=(-12u^{2}+48u^{3}+\cdots\ )\ du$
Hence, taking
$Z_{1}$ $=4x+8iy+12t+\phi_{1}$
(10) $Z_{2}$ $=-12t+\phi_{2}$
$Z_{3}$ $=4x-8iy+12t+\phi_{3}$
$Z_{4}$ $=-12t+\phi_{4}$
and substituting into (8), we obtain what is essentially a $\tau$-function for KP1 solution (note that the complex rescaling of $y$ to $i\cdot y$ has been incorporated here). The $\phi_{j}$ are arbitrary constant parameters (analogous to “phase factors” for soliton solutions). They must be chosen appropriately to obtain a regular real solution. In fact, for “most” values of the $\phi_{j}$ the resulting KP1 solutions produced by this recipe will still be non-regular, and they will also take non-real values at some real $(x,y,t)$. We will see how to overcome these difficulties in the next section.
5. Finding a real regular solution
We begin with a simple observation related to the form of (8) and (4). As long as $x,t$ are taken to be real, the only possible terms contributing non-real values in the substituted theta-function are those containing odd powers of $y$–that is, $y,y^{3}$, and $y^{5}$–together with arbitrary powers of $x,t$ giving a total degree at most $6$. The coefficient of $y^{5}$ yields the terms
$(-24\phi_{1}+24\phi_{3}+384)i$
$(44-2\phi_{4}+2\phi_{2})i.$
$+(1125/2)t+(249/4)x^{4}+(19521/4)t^{4}.$
Theorem 4.
Proof.
That the $u(x,y,t)$ produced by (2) takes only real values for real $(x,y,t)$ is a consequence of the determination of the $\phi_{j}$ described above and is also clearly visible from the form of (11). To show that this is a regular solution (a “multi-lump”) we will show that the polynomial in (11) is a sum of squares of real polynomials with a nonzero constant term. This will complete the proof. To begin, we note that the polynomial analog of the theta-function from (8) above can actually be rewritten in the form
$+36(Z_{1}Z_{3}+10Z_{1}-6Z_{3}-56)(Z_{1}Z_{3}+6Z_{1}-10Z_{3}-56).$
Part of this reflects the general patterns determined in the proof of Theorem 5.3 of [2]. The two factors of degree 3 on the first line are actually the implicit equations of certain translates of the theta-divisors for the two partial normalizations of the bicuspidal curve, where one of the cusps is smoothed and the other one remains untouched. The other term is of lower total degree 4 and it happens to factor in this way in this case. When the values from (4) with $\phi_{j}$ as above are substituted into this polynomial, the two factors of total degree 3 in the term on the first line become complex conjugates, and the product has the form $(A+iB)(A-iB)=A^{2}+B^{2}$, where $A,B$ are polynomials with real coefficients. The factors on the second line yield
$36((4x+12t+10)^{2}+64y^{2}+4)((4x+12t+6)^{2}+64y^{2}+4)$
Figure 1. The solution (11) at $t=-1.375$.
Figure 2. The solution (11) at $t=-0.9375$.
Figure 3. The solution (11) at $t=-0.4375$.